On Tuesday night, when I added the Si spacer to the clamp, I measured a Q of ~46000, but I noted that it had been increasing up to that point, likely due to the residual gas damping (see CRYO:1214). Last night, I made another measurement and found it to be much higher, at ~1.2 x 105 (tau ~ 350 s). This is much better than we have seen at room temperature thus far, so it looks like my spacer addition has helped.
I remeasured this an hour or so later and saw no appreciable increase. I checked again today and it appears as though it may have increased slightly, but it was hard to say for sure due to higher environmental noise. Really, we need the steady-state ringdown to make a good measurement at this level.
The LN2 dewar was refilled today, so I filled the cryostat and we'll see how it looks at low temperature tomorrow.